Determine the y-intercept, the zeros, the number of turning points, and the behavior of the graph and sketch the graph of the following polynomial functions.



1. P(x)=-2x³-22x²-60x


2. P(x)=6x³+5x²-2x-1


4. P(x)=x⁴+x³-10x²+8




1
Expert's answer
2021-11-21T11:59:55-0500

1. "P(x)=-2x^3-22x^2-60x"

"y" -intercept: "x=0"

"P(0)=-2(0)^3-22(0)^2-60(0)=0"

Point "(0,0)"

The graph passes through the origin.

zeros: "y=0=>-2x^3-22x^2-60x=0"


"-2x(x^2+11x+30)=0"

"-2x(x+6)(x+5)=0"

"zeros; -6, -5, 0"

Point "(-6,0)", Point "(-5,0)", Point "(0,0)"


"y'=-6x^2-44x-60"

"y'=0=>-6x^2-44x-60=0"

"3x^2+22x+30=0"

"D=(22)^2-4(3)(30)=124"

"x=\dfrac{-22\pm\sqrt{124}}{2(3)}=\dfrac{-11\pm\sqrt{31}}{3}"

There are two turning points.

"y\to \infin" as "x\to-\infin"

"y\to -\infin" as "x\to\infin"



2. "P(x)=6x^3+5x^2-2x-1"

"y" -intercept: "x=0"

"P(0)=6(0)^3+5(0)^2-2(0)-1=-1"

Point "(0,-1)"


zeros: "y=0=>6x^3+5x^2-2x-1=0"


"6x^3+6x^2-x^2-x-x-1=0"

"6x^2(x+1)-x(x+1)-(x+1)=0"




"(x+1)(6x^2-x-1)=0"

"(x+1)(2x-1)(3x+1)=0"


"zeros; -1, -1\/3, 1\/2"

Point "(-1,0)", Point "(-1\/3,0)", Point "(1\/2,0)"


"y'=18x^2+10x-2"

"y'=0=>18x^2+10x-2=0"

"9x^2+5x-1=0"

"D=(5)^2-4(9)(-1)=61"

"x=\dfrac{-5\pm\sqrt{61}}{2(9)}"

There are two turning points.

"y\to -\infin" as "x\to-\infin"

"y\to \infin" as "x\to\infin"




4. "P(x)=x^4+x^3-10x^2+8"

"y" ​-intercept:"x=0"

"P(0)=(0)^4+(0)^3-10(0)^2+8=8"

Point "(0,8)"


zeros: "y=0=>x^4+x^3-10x^2+8=0"


"x^4-x^3+2x^3-2x^2-8x^2+8=0"

"x^3(x-1)+2x^2(x-1)-8(x+1)(x-1)=0"




"(x-1)(x^3+2x^2-8x-8)=0"

"x_1=1"

"P(-4)=(-4)^4+(-4)^3-10(-4)^2+8=40>0"


"P(-3)=(-3)^4+(-3)^3-10(-3)^2+8=-28<0"

There is "c_1\in(-4, -3)" such that "y(c_1)=0."


"P(-1)=(-1)^4+(-1)^3-10(-1)^2+8=-2<0"


"P(0)=(0)^4+(0)^3-10(0)^2+8=8>0"

There is "c_2\in(-1, 0)" such that "y(c_2)=0."



"P(2)=(2)^4+(2)^3-10(2)^2+8=-8<0"


"P(3)=(3)^4+(3)^3-10(3)^2+8=26>0"

There is "c_3\in(2, 3)" such that "y(c_3)=0."

"zeros; c_1, c_2, 1, c_3."

Point "(-1,0)", Point "(-1\/3,0)", Point "(1\/2,0)"


"y'=4x^3+3x^2-20x"

"y'=0=>4x^3+3x^2-20x=0"

"x(4x^2+3x-20)=0"

"x_1=0"

Or


"4x^2+3x-20=0"

"D=(3)^2-4(4)(-20)=329"

"x=\dfrac{-3\pm\sqrt{329}}{2(4)}"

There are three turning points.

"y\to\infin" ​as "x\to-\infin"

"y\to \infin" ​as "x\to\infin"